# PROPS: Progressively Private Self-alignment of Large Language Models

### v1

@@ 3.1 Theoretical Results for PROPS
In this section, we analyze the Sub-optimality gap, which captures the gap between the optimal non-private DPO policy parameters $\theta^{*}$, and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$. To obtain a bound on the Sub-optimality gap, we follow similar assumptions as those made in Chowdhury et al. (2024a), including expressing the model as a log-linear policy and assuming smoothness by placing bounds on the policy and its gradients. This gap is formalized in Theorem 1; we provide more details on the assumptions and derivation in Section A.5.
@@
Inputs: Fine-tuned Model $M_{0}$, Dataset $\mathcal{D}$, Stages $K$, Privacy parameters $(\epsilon,\delta)$
Output: Aligned model $M_{K}$
 Perform RR on $\mathcal{D}$, such that $\mathcal{D}\xrightarrow{RR(\gamma_{\epsilon})}\mathcal{D}^{\prime}$
 Partition the dataset $\mathcal{D}^{\prime}$ into $K$-disjoint datasets such that $\mathcal{D}^{\prime}=D_{1}\cup D_{2}\cup\ldots\cup D_{K}$
 Flip labels using Randomized Response (RR) with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$
 Align model $M_{0}$ with dataset $D_{1}$ to obtain $M_{1}$
for $k=2,3,\ldots,K$ do
  Generate labels $\ell_{M_{k-1}}^{D_{k}}$ for dataset $D_{k}$ using the model $M_{k-1}$ (from previous stage)
  Obtain $\ell_{RR}^{D_{k}}$ and $\ell_{M_{k-1}}^{D_{k}}$ and obtain the maximum likelihood estimator (MLE) $\Lambda$ according to equation 9 and generate label as:
$\small\ell_{\text{PROPS}}^{D_{k}}=\begin{cases}1,\quad\text{ If }\Lambda\leq 0\\ 0,\quad\text{ If }\Lambda>0.\end{cases}$
  Align model $M_{k-1}$ on dataset $D_{k}$ with PROPS labels ($\ell_{\text{PROPS}}^{D_{k}}$) to get model $M_{k}$
end for
Output: Aligned model $M_{K}$ on dataset $\mathcal{D}$
@@
Remark 2 Distinction from Label-DP, Multi-Stage RR and PATE: While the notion of ($\epsilon,\delta$)-Preference Privacy is motivated from the notion of $(\epsilon,\delta)$-Label DP (Chaudhuri & Hsu, 2011; Ghazi et al., 2021a), there are distinctions between the two frameworks. In Label-DP, only labels are treated as private, and noise is added in proportion to the privacy budget to preserve the privacy of the dataset’s labels. In addition, RR (Warner, 1965) is a direct approach to implement preference privacy, as the preference $\ell_{i}$ of a data entry is flipped with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$. While PROPS shares similarities with Multi-Stage RR Ghazi et al. (2021b) and PATE Papernot et al. (2018), it differs significantly in approach and application. Unlike Multi-Stage RR, which relies on simple sampling for combining noisy labels and model predictions, PROPS uses an MLE-based approach for principled integration. PATE protects privacy for all features using parallel training, whereas PROPS targets preference privacy with a sequential, iterative approach that improves alignment by building on earlier models while preserving privacy.
@@ 4 Experiments and Discussion
Datasets and models: In our experiments and validation, we have used (1) three datasets (jondurbin/truthy-dpo-v0.1, Anthropic HH-RLHF, and AlpacaEval) and (2) three different models of varying sizes: Pythia-1B EleutherAI (2024) , GPT2-Large (774M) OpenAI (2024a) and GPT2-Medium (355M) OpenAI (2024b). We have adopted a similar experimental setup as the prior works Rafailov et al. (2024); Chakraborty et al. (2024); von Werra et al. (2020)Lior-Baruch (2024). The code for PROPS is publicly available [^1] https://anonymous.4open.science/r/PROPS-2025.
@@
PROPS vs DPSGD: In Table 2, we present Win-Tie rates comparing our proposed algorithm PROPS with the conventional DP-SGD algorithm for GPT2-Large and GPT2-Medium models on the truthy-dpo-v0.1 dataset. DP-SGD was ran using the Gaussian mechanism for 1 epoch with $\delta=10^{-10}$, a gradient clipping threshold of 10, and a batch size of 4 for 2 epochs. As the results indicate, PROPS is able to consistently outperform DP-SGD at higher privacy regimes ($\epsilon=0.1,0.5,1$) for both models. This indicates that while DP-SGD attempts to additionally protect the prompts and responses, it suffers a significant drop in utility for smaller privacy budgets. Additional results are presented in Section A.6. One critical distinction to highlight is that PROPS ensures a pure DP guarantee (i.e. $(\epsilon,0)$-DP) while DP-SGD provides an approximate DP guarantee, denoted as $(\epsilon,\delta)$-DP. We present results for PROPS vs DP-SGD on HH-RLHF, AlpacaEval and truthy-dpo-v0.1 for three privacy parameters. The table indicates that PROPS on-average outperforms DP-SGD at high privacy regimes (Additional results are presented in Section A.6).
@@
Results on multi-stage PROPS: We now report results to observe if performance gains can be attained by increasing the number of stages used in PROPS. In Table 3, we present results comparing 2-stage and 3-stage PROPS using the truthy-dpo-v0.1 dataset with GPT2-Large. Results indicate that across high privacy budgets, 2-stage PROPS outperforms 3-stage PROPS. This implies that further hyperparameter tuning under privacy constraints is required for 3 or more stages. We leave this as future work as our results up to this point have indicated that the 2-stage results effectively illustrate the key performance trends under varying privacy constraints.
@@ 5 Conclusions
This work aims to advance the field of privacy-preserving machine learning by addressing the challenge of protecting human labelers’ preferences during the alignment of Large Language Models (LLMs). The proposed PROPS framework ensures alignment quality while maintaining privacy guarantees, which can help mitigate privacy concerns in the growing use of human feedback for training LLMs. By improving the privacy-utility tradeoff, this approach supports the use of such language models while fostering trust in systems that rely on human data. Due to resource constraints, we have used “smaller” language models (e.g. GPT2-Medium, GPT2-Large, Pythia-1B); however, our results still indicate the effectiveness of PROPS in ensuring preference level privacy. Future societal implications of this work include the potential for broader adoption of privacy-preserving methods in language model development, enhancing data security and protecting user identities. However, as there is a need to monitor and ensure that such systems are deployed responsibly and in alignment with ethical guidelines. This work contributes positively to the field by promoting ethical considerations in LLM training and alignment.
@@ A.1 Training Details & Comparison of Complexity
truthy-dpo-v0.1: For this dataset, 15% of the data was used for SFT. The remaining 75% of the data was designated for DPO training. This 75% segment was divided into two halves, with three epochs of DPO run on each half. Subsequently, the dataset was filtered to include only preference pairs generated by prompting a Large Language Model (LLM) to act as ”an honest and helpful assistant.” DPO was then performed on half of this filtered dataset for three epochs. Win-Tie-Loss rates were calculated using the remaining 10% of the Truthy-DPO-v0.1 dataset, which consists of 100 prompts.
@@ A.3 Estimator for $\gamma_{M_{1}}$: Proof of (8)
Validation of our estimation procedure for $\gamma^{*}_{M_{1}}$: To validate the robustness of our estimation process for obtaining $\hat{\gamma_{M_{1}}}$, we compare the estimated error rate of model $M_{1}$ with its true counterpart (the ”oracle” error rate of $\gamma_{M_{1}}^{*}$). Specifically, Table 5 presents a comparison between the estimated error rate $\hat{\gamma}_{M_{1}}$, computed using equation 8 from the paper, and the oracle error rate $\hat{\gamma}_{M_{1}}^{*}$, obtained by evaluating $M_{1}$’s predictions on the unperturbed preference data against the ground truth (i.e., the original preferences prior to flipping). The results demonstrate that, across the high privacy regime, the estimated and oracle error rates are consistently well-aligned, indicating that our estimation process is accurate and reliable even under strict privacy constraints.
Table 5: Comparable estimations on flipping for our method ($\hat{\gamma_{M_{1}}}$) compared to the model flipping probability obtained from the ”oracle” setting $\gamma_{M_{1}}^{*}$.
@@ A.4 Proof Sketch of Theorem 1: Sub-optimality gap for PROPS
Chowdhury et al. (2024a) provides a bound on the sub-optimality gap between the rewards obtained using an optimal model aligned under DPO with noiseless preference data and a model aligned under their proposed robust DPO (rDPO) method, which accounts for noisy (i.e. flipped) preference data. Their result characterizes how many preference samples are needed at different noise levels to ensure that the loss in rewards (relative to the original DPO method) does not exceed a certain bound. Chowdhury et al. (2024a) assume that the model can be expressed as a log-linear policy (i.e. as a function of a feature map $\phi(x,y)^{T}$ and the weights of the last layer $\theta$). They assume that the policy (characterized by the parameters $\theta$) and its first and second order gradients are bounded, to provide bounds and Lipschitz guarantees on the difference between rewards attained for the chosen and rejected responses. It is also assumed that the parameters $\theta$ are in the following set: $\{\theta\in\mathbb{R}^{d}|\sum_{i=1}^{d}\theta_{i}=0\}$. Additionally, they assume that the fine-tuned model (i.e. model before alignment) has a good coverage of the feature space, to ensure that the relative condition number $\kappa$ between the covariance matrices of the aligned and fine-tuned policies is small. They prove that their sub-optimality gap is as follows:
@@
In each of the four possible scenarios, we observe that for $\gamma_{M_{1}}<\gamma_{\epsilon}$, the predictions of the MLE match the predictions of $M_{1}$. This implies that $\gamma_{\text{MLE}}=\gamma_{M_{1}}$. Conversely, when $\gamma_{\epsilon}<\gamma_{M_{1}}$, performing the same analysis shows that the MLE will match the predictions of RR, resulting in $\gamma_{\text{MLE}}=\gamma_{\epsilon}$. Therefore, the flipping rate of PROPS can be described as $\gamma_{\text{MLE}}=\min(\gamma_{\epsilon},\gamma_{M_{1}})$. Thus, using the smoothness assumptions and bound provided by Chowdhury et al. (2024a), the sub-optimality gap between between the optimal non-private DPO policy parameters $\theta^{*}$ and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$ can be obtained as:

### v2

@@ 3.1 Theoretical Results for PROPS
In this section, we analyze the Sub-optimality gap, which captures the gap between the optimal non-private DPO policy parameters $\theta^{*}$, and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$. To obtain a bound on the Sub-optimality gap, we follow similar assumptions as those made in chowdhury2024provably, including expressing the model as a log-linear policy and assuming smoothness by placing bounds on the policy and its gradients. This gap is formalized in Theorem 1; we provide more details on the assumptions and derivation in Section A.5.
@@
1:Fine-tuned model $M_{0}$, dataset $\mathcal{D}$, number of stages $K$, privacy parameters $(\epsilon,\delta)$
2:Aligned model $M_{K}$
3:Perform RR on $\mathcal{D}$ to obtain $\mathcal{D}^{\prime}$ $\triangleright$ $\mathcal{D}\xrightarrow{RR(\gamma_{\epsilon})}\mathcal{D}^{\prime}$
4:Partition $\mathcal{D}^{\prime}$ into $K$ disjoint subsets: $\mathcal{D}^{\prime}=D_{1}\cup D_{2}\cup\dots\cup D_{K}$
5:Set flip probability $\gamma_{\epsilon}=\dfrac{1}{1+e^{\epsilon}}$
6:Align $M_{0}$ on $D_{1}$ (using $R$ labels) to obtain $M_{1}$
7:for $k=2$ to $K$ do
8:  Use $M_{k-1}$ to generate labels $\ell_{M_{k-1}}^{D_{k}}$ on $D_{k}$
9:  Obtain $\ell_{\mathrm{RR}}^{D_{k}}$ and $\ell^{D_{k}}_{M_{k-1}}$ for dataset $D_{k}$ and obtain Maximum Likelihood Estimator (MLE) $\Lambda$ according to Equation 5, and generate labels as $\ell_{\text{PROPS}}^{D_{k}}=\begin{cases}1,&\text{if }\Lambda\leq 0,\\ 0,&\text{if }\Lambda>0,\end{cases}$
10:  Align $M_{k-1}$ on $D_{k}$ with PROPS labels $\ell_{\text{PROPS}}^{D_{k}}$ to obtain $M_{k}$
11:end for
12:return $M_{K}$
@@
Remark 2 Distinction from Label-DP, Multi-Stage RR and PATE: While the notion of ($\epsilon,\delta$)-Preference Privacy is motivated from the notion of $(\epsilon,\delta)$-Label DP (chaudhuri2011sample; NEURIPS2021_e3a54649), there are distinctions between the two frameworks. In Label-DP, only labels are treated as private, and noise is added in proportion to the privacy budget to preserve the privacy of the dataset’s labels. In addition, RR (warner1965randomized) is a direct approach to implement preference privacy, as the preference $\ell_{i}$ of a data entry is flipped with probability $\gamma_{\epsilon}=\frac{1}{1+e^{\epsilon}}$. While PROPS shares similarities with Multi-Stage RR ghazi2021deep and PATE papernot2018scalable, it differs significantly in approach and application. Unlike Multi-Stage RR, which relies on simple sampling for combining noisy labels and model predictions, PROPS uses an MLE-based approach for principled integration. PATE protects privacy for all features using parallel training, whereas PROPS targets preference privacy with a sequential, iterative approach that improves alignment by building on earlier models while preserving privacy.
@@ 4 Experiments and Discussion
Datasets and models: In our experiments and validation, we have used (1) three datasets (jondurbin/truthy-dpo-v0.1, Anthropic HH-RLHF, and AlpacaEval) and (2) three different models of varying sizes: Pythia-1B Pythia-1b , GPT2-Large (774M) GPT2-Large and GPT2-Medium (355M) GPT2-Medium. We have adopted a similar experimental setup as the prior works rafailov2024directpreferenceoptimizationlanguage; chakraborty2024transferqstarprincipled; vonwerra2022trllior-baruch. The code for PROPS is publicly available [^1] https://anonymous.4open.science/r/PROPS-2025.
@@
PROPS vs DPSGD: In Table 2, we present Win-Tie rates comparing our proposed algorithm PROPS with the conventional DP-SGD algorithm for GPT2-Large and GPT2-Medium models on the truthy-dpo-v0.1 dataset. DP-SGD was ran using the Gaussian mechanism for 1 epoch. We set a gradient clipping threshold C = 10, as a lower value (e.g., C=1) introduced significant clipping bias that harmed utility. Our choice of $\delta=10^{-10}$ is intentionally conservative and stricter than the common $1/N$ heuristic to establish a strong privacy baseline.
As the results indicate, PROPS is able to consistently outperform DP-SGD at higher privacy regimes ($\epsilon=0.1,0.5,1$) for both models. This indicates that while DP-SGD attempts to additionally protect the prompts and responses, it suffers a significant drop in utility for smaller privacy budgets. Additional results are presented in Section A.6. One critical distinction to highlight is that PROPS ensures a pure DP guarantee (i.e. $(\epsilon,0)$-DP) while DP-SGD provides an approximate DP guarantee, denoted as $(\epsilon,\delta)$-DP. We present results for PROPS vs DP-SGD on HH-RLHF, AlpacaEval and truthy-dpo-v0.1 for three privacy parameters. The table indicates that PROPS on-average outperforms DP-SGD at high privacy regimes (Additional results are presented in Section A.6).
@@
Results on multi-stage PROPS:
We now report results comparing 2-stage and 3-stage PROPS using the truthy-dpo-v0.1 dataset with GPT2-Large in Table 3. Our findings highlight an interesting tradeoff in the PROPS pipeline: the optimal number of alignment stages depend on the available privacy budget.
When the privacy budget is low (small $\epsilon$), each human-labeled preference needs to be heavily perturbed, resulting in high label noise. Under this regime, increasing the number of stages provides diminishing returns: early-stage models are themselves noisy, and relabeling through additional stages can further propagate this noise. Moreover, since the dataset must be partitioned across stages, later stages have less effective data to counteract the noise. As a result, a smaller number of stages (e.g., $K=2$) tends to yield better alignment in high-privacy settings.
In contrast, when the privacy budget is moderate or higher (larger $\epsilon$), each preference label is less noisy, and earlier-stage models produce more reliable pseudo-labels. This improves the quality of relabeling in subsequent stages, allowing additional stages to refine alignment rather than amplify errors. Consequently, a larger number of stages (e.g., $K=3$) can begin to provide measurable gains, as evidenced in Table 3, where at $\epsilon=2.0$ the 3-stage setup achieves $44.0\%$ wins compared to $38.4\%$ for 2 stages.
These results suggest that number of stages in PROPS should scale with privacy budget and data availability: fewer stages are more robust under strong privacy constraints and limited samples, whereas additional stages can be beneficial once the signal-to-noise ratio of the preference data improves.
@@ 5 Conclusions
This work aims to advance the field of privacy-preserving machine learning by addressing the challenge of protecting human labelers’ preferences during the alignment of Large Language Models (LLMs). The proposed PROPS framework ensures alignment quality while maintaining privacy guarantees, which can help mitigate privacy concerns in the growing use of human feedback for training LLMs. By improving the privacy-utility tradeoff, this approach supports the use of such language models while fostering trust in systems that rely on human data. Due to resource constraints, we have used “smaller" language models (e.g. GPT2-Medium, GPT2-Large, Pythia-1B); however, our results still indicate the effectiveness of PROPS in ensuring preference level privacy. Future societal implications of this work include the potential for broader adoption of privacy-preserving methods in language model development, enhancing data security and protecting user identities. However, as there is a need to monitor and ensure that such systems are deployed responsibly and in alignment with ethical guidelines. This work contributes positively to the field by promoting ethical considerations in LLM training and alignment.
5.1 Acknowledgments
This work was supported by NIH Award R01-CA261457-01A1 and also by the US Department of Energy, Office of Science, Office of Advanced Scientific Computing under Award Number DE-SC-ERKJ422, and US NSF under Grants CCF-2100013, CNS-2209951, CNS-2317192.
@@ A.1 Training Details & Comparison of Complexity
truthy-dpo-v0.1: For this dataset, 15% of the data was used for SFT. The remaining 75% of the data was designated for DPO training. This 75% segment was divided into two halves, with three epochs of DPO run on each half. Subsequently, the dataset was filtered to include only preference pairs generated by prompting a Large Language Model (LLM) to act as "an honest and helpful assistant." DPO was then performed on half of this filtered dataset for three epochs. Win-Tie-Loss rates were calculated using the remaining 10% of the Truthy-DPO-v0.1 dataset, which consists of 100 prompts.
@@
Comparison of Wall-Clock Time : While Big-O analysis provides theoretical insights into computational complexity, empirical comparisons of wall-clock time are crucial for understanding the practical overhead introduced by PROPS. We benchmarked a standard DPO run against our two-stage PROPS procedure using GPT-2 Medium and GPT-2 Large models, with results summarized in Table 5. For a given model and dataset size, the core DPO training times are nearly identical across methods. The primary additional cost in PROPS arises from Stage 2, where pseudo labels ($l_{M_{1}}$) are generated on the second data partition ($D_{2}$). In our experiments, this pseudo-label generation constitutes the dominant component of the overhead. Nevertheless, this extra computation represents a necessary and well-justified trade-off. As shown by the win-rate results in Section 4 (e.g., Tables 1) and Appendix A.6 (Tables 7), the inclusion of this step enables PROPS to achieve markedly higher model utility and alignment quality compared to baselines such as RR and DP-SGD—particularly under stringent privacy constraints ($\epsilon\leq 1$) where these alternatives exhibit substantial performance degradation. In essence, the additional compute directly translates into meaningful improvements in model performance and alignment.
Table 5: Wall-clock time comparison (500 entries)
Model Method DPO Training Label Generation Total Time
GPT2-M $R$ 42 s 0 s 42 s
GPT2-M PROPS 42 s 4 min 31 s 5 min 13 s
GPT2-L $R$ 1 min 53 s 0 s 1 min 53 s
GPT2-L PROPS 1 min 53 s 11 min 36 s 13 min 29 s
While PROPS incurs additional computational cost due to the label generation step, this overhead is justified by its substantially improved alignment quality, particularly in high-privacy settings. As demonstrated in our main experimental results, PROPS achieves significantly higher win rates compared to RR and DP-SGD under stringent privacy budgets ($\epsilon\leq 1$), indicating that the extra computation translates directly into better model utility where conventional methods falter.
@@ A.3 Estimator for $\gamma_{M_{1}}$: Proof of (8)
Validation of our estimation procedure for $\gamma^{*}_{M_{1}}$: To validate the robustness of our estimation process for obtaining $\hat{\gamma_{M_{1}}}$, we compare the estimated error rate of model $M_{1}$ with its true counterpart (the "oracle" error rate of $\gamma_{M_{1}}^{*}$). Specifically, Table 6 presents a comparison between the estimated error rate $\hat{\gamma}_{M_{1}}$, computed using equation 8 from the paper, and the oracle error rate $\hat{\gamma}_{M_{1}}^{*}$, obtained by evaluating $M_{1}$’s predictions on the unperturbed preference data against the ground truth (i.e., the original preferences prior to flipping). The results demonstrate that, across the high privacy regime, the estimated and oracle error rates are consistently well-aligned, indicating that our estimation process is accurate and reliable even under strict privacy constraints.
Table 6: Comparable estimations on flipping for our method ($\hat{\gamma_{M_{1}}}$) compared to the model flipping probability obtained from the "oracle" setting $\gamma_{M_{1}}^{*}$.
@@ A.4 Proof Sketch of Theorem 1: Sub-optimality gap for PROPS
chowdhury2024provably provides a bound on the sub-optimality gap between the rewards obtained using an optimal model aligned under DPO with noiseless preference data and a model aligned under their proposed robust DPO (rDPO) method, which accounts for noisy (i.e. flipped) preference data. Their result characterizes how many preference samples are needed at different noise levels to ensure that the loss in rewards (relative to the original DPO method) does not exceed a certain bound. chowdhury2024provably assume that the model can be expressed as a log-linear policy (i.e. as a function of a feature map $\phi(x,y)^{T}$ and the weights of the last layer $\theta$). They assume that the policy (characterized by the parameters $\theta$) and its first and second order gradients are bounded, to provide bounds and Lipschitz guarantees on the difference between rewards attained for the chosen and rejected responses. It is also assumed that the parameters $\theta$ are in the following set: $\{\theta\in\mathbb{R}^{d}|\sum_{i=1}^{d}\theta_{i}=0\}$. Additionally, they assume that the fine-tuned model (i.e. model before alignment) has a good coverage of the feature space, to ensure that the relative condition number $\kappa$ between the covariance matrices of the aligned and fine-tuned policies is small. They prove that their sub-optimality gap is as follows:
@@
In each of the four possible scenarios, we observe that for $\gamma_{M_{1}}<\gamma_{\epsilon}$, the predictions of the MLE match the predictions of $M_{1}$. This implies that $\gamma_{\text{MLE}}=\gamma_{M_{1}}$. Conversely, when $\gamma_{\epsilon}<\gamma_{M_{1}}$, performing the same analysis shows that the MLE will match the predictions of RR, resulting in $\gamma_{\text{MLE}}=\gamma_{\epsilon}$. Therefore, the flipping rate of PROPS can be described as $\gamma_{\text{MLE}}=\min(\gamma_{\epsilon},\gamma_{M_{1}})$. Thus, using the smoothness assumptions and bound provided by chowdhury2024provably, the sub-optimality gap between between the optimal non-private DPO policy parameters $\theta^{*}$ and the policy parameters obtained through two-stage PROPS $\hat{\theta}_{\text{PROPS}}$ can be obtained as:
@@
Limitations of Theoretical Analysis: Our theoretical analysis relies on idealized assumptions—such as log-linear policies, smoothness, and feature coverage—that enable tractable mathematical treatment but may not hold for modern Transformer-based architectures. These simplifying assumptions are standard in the literature and are intended to provide interpretable intuition about the underlying mechanisms rather than exact quantitative guarantees. In particular, the theory correctly predicts how performance scales with key parameters such as the effective noise rate of the first-stage MLE ($\min(\gamma_{M_{1}},\gamma_{\epsilon})$) and the second-stage sample size ($n_{2}$), offering a principled explanation of the observed trade-offs. As a result, while the derived scaling trends offer valuable insight and a principled sanity check on empirical behavior, the bounds should not be interpreted as directly predictive of large-scale model performance. Bridging this gap between theoretical idealizations and real-world LLM architectures remains an important direction for future work.